Two Zuc point changes are located on the x-axy. One is at x= Im and the other is at n=-1m.





(a). Determine the electric field on the y-axis at





y=0.5m

1
Expert's answer
2022-03-26T15:31:44-0400

"E=2\frac{ke}{\sqrt{x^2+y^2}}\frac{\sqrt 5}5=1.2~\frac{nV}m ."